The sum of the ages of Akin and Dop is 35years. The sum of twice Akins age and three times Dop’s age is
89years. Find their present ages.

Found 2 solutions by greenestamps, Cromlix:
Answer by greenestamps(13200)



(1) The sum of the ages of Akin and Dop is 35 years [given]

(2) The sum of twice Akin's age and three times Dop's age is 89 [given]

(3) The sum of twice Akin's age and twice Dop's age is 2*35 = 70 years [from (1)]

(4) Dop's age is 19 [comparing (2) and (3)]

(5) Akin's age is 16 [comparing (1) and (4)]

----------------------------------------------

In formal algebraic terms....

(1) A + D = 35 [given]
(2) 2A+3D = 89 [given]
(3) 2A+2D = 70 [multiply (1) by 2]
(4) D = 19 [subtract (3) from (2)]
(5) A = 16 [subtract (4) from (1)]


Answer by Cromlix(4381)

Hi there,
A + D = 35..........(1)
2A + 3D = 89........(2)
Multiply (1) by 2
2A + 2D = 70........(1)
2A + 3D = 89........(2)
Subtract (1) from (2)
D = 19
Substitute D = 19 into (1)
A + D = 35..........(1)
A + 19 = 35
A = 35 - 19
A = 16
Akin is 16 yrs old
Dop is 19 yrs old.
Hope this helps :-)